# Gradient-free training of recurrent neural networks

### v1

@@ 4 Computational experiments
We now discuss a series of experiments designed to illustrate the benefits and challenges of our construction approach. We compare our method to the state-of-the-art iterative gradient-based method called shPLRNN, which we explain in Appendix D. For the real-world weather dataset, we also compare our approach with a long short-term memory (LSTM) model. Furthermore, due to the similarities our method bears with reservoir models, we also compare with an established reservoir model, namely an echo state network (ESN), further explained in Appendix E. Details on the datasets can be found in Appendix F, hyperparameters for all models are given in Appendix H, the evaluation metrics are explained in Appendix G and a further comparison discussion as well as the hardware details are provided in Appendix H. The code to reproduce the experiments from the paper, as well as an up-to-date code base, can be found at
@@
https://gitlab.com/felix.dietrich/swimrnn
In Table 1, we list the quantitative results of the experiments without control. Each entry stems from five different runs, where the random seed is changed in order to ensure a more robust result. We give the mean over these five runs, as well as the minimum and maximum among them.
@@ 4.1 Simple ODEs: Van der Pol Oscillator
We consider the Van der Pol oscillator system for a simple illustration of our method. A sampled RNN with a $tanh$ activation and a single hidden layer of width 80 is used, and the prediction method follows Algorithm 3. The model is evaluated on test data; the averaged error and training time are reported in Table 1. One trajectory from the test set is visualized in Figure 2. It should be noted that predictions start with an initial condition from the test dataset which is used to make the first prediction, and afterwards continue using this prediction as an input to predict the next state, without information from the ground-truth dataset. In the results we observe a very stable trajectory over a long prediction horizon, and furthermore all eigenvalues of the Koopman operator are inside the unit disk (see Appendix C), thus we are certain that the model is stable. This experiment is also significant due to the periodic nature of the system, which is captured with our model, although neural network architectures in general struggle to capture periodicity (Ziyin et al., 2020). Compared to the gradient-descent trained shPLRNN our method is much faster and achieves higher forecasting accuracy, with lower MSE as prediction error. Compared with an ESN, our model has a shorter fit time and a smaller error. In addition, the hyperparameter search is much simpler and less time consuming for our method since there are fewer hyperparameters to tune.
@@ 4.4 Example with control inputs: Forced Van der Pol Oscillator
We consider again the Van der Pol oscillator, where now the second coordinate, $h_{2}$, is controlled with an external input $x$, and using a sampled RNN model we perform model predictive control (MPC) as in Algorithm 4. We let $\mathcal{G}_{\hat{M}}$ be the identity function, and let $B$ map from $\mathcal{X}$ to $\mathbb{R}^{M}$ (adding nonlinearity for $x$ did not yield different results for this system, and is described in Section H.1.1). The sampled recurrent neural network is then the identity $\mathcal{G}_{\hat{M}}$ and $\mathcal{F}_{M}$ is a hidden layer of width 128 and $tanh$ activation. This network is then passed as a surrogate model to a linear-quadratic regulator (LQR). The network, in combination with the LQR, can successfully steer the state to the target state (see Figure 5). We consider five different runs, where only the random seed is varied, and obtain the mean controller cost to be 125.92 and the mean training time of 1.122 seconds. The $L^{2}$ norm of the state is also tracked over time, for five different runs we show the norms and the pointwise mean (over the runs) in Figure 5. This experiment highlights a key advantage of our model, which allows for modelling a nonlinear system such as the Van der Pol oscillator using a linear controller such as LQR. This implies that the well-established tools from linear control theory can be applied to non-linear systems using our method.
@@ 4.5 Example with real-world data
We apply our approach to the climate data presented in TensorFlow (2024). The dataset contains a time series of 14 weather parameters recorded in Jena (Germany) between January 1st, 2009, and December 31st, 2016. We used the largest consecutive time range for the experiments and trained a sampled RNN to predict the temperature parameter (Figure 6). We also trained an autoregressive LSTM model described in TensorFlow (2024) and a shPLRNN as iterative baselines.
The data offers freedom in choosing the sizes of the time delay and prediction horizons. We decided to fix the time delay to one week and set two separate experiments with a prediction horizon of one day and one week. In the sampled RNN and LSTM experiments, we performed a grid search for each model with hyperparameters specified in Section H.4.1.
Table 1 shows the averaged training time and error metrics for the two selected horizons (day and week). We observe that the sampled and iterative models perform similarly in the case of a shorter horizon, while sampling offers much faster training. When considering a one-week horizon, shPLRNN predictably outperforms LSTM, but sampled RNN is still orders of magnitude faster. We also note the prediction horizon does not influence the training time of the sampled model that agrees with the Algorithm 2. When comparing predictions for the longer horizon in Figure 6, we notice that the LSTM struggles to predict the high-frequency fluctuations of the measurement, but the sampled RNN and shPLRNN successfully capture them. Figure 6 also highlights the deficiency of the MSE metric because a low mean error does not always correspond to accurate predictions, as illustrated by all three models. Overall, we conclude that sampled RNNs can successfully capture chaotic real-world dynamics and produce results comparable to the iterative models while offering a significant speed-up in training.
@@ 5 Conclusion
Benefits of the approach In many examples, we demonstrate that we can train accurate recurrent networks orders of magnitude faster compared to networks trained by iterative methods. We also observe that the training approach works with a very small amount of training data. The direct connection to Koopman operator theory allows us to draw on existing theoretical results for dynamical systems, which we use to prove convergence in the limit of infinite width. In addition, the same connection allows us to apply linear control theory to non-linear systems.
@@
Remaining challenges and future work Recurrent networks are often used for tasks in computer vision and natural language processing. These tasks require network architectures beyond feed-forward networks, like convolutional neural networks or transformers. The sampling scheme we use to construct the hidden weights of the neurons is currently not useful in constructing parameters for such architectures, but it is an intriguing challenge to work towards sampling them. Remaining challenges in the theoretical work include extending the theory shown in this paper to controlled systems, as well as bridging the Koopman theory for continuous dynamical systems and NeuralODEs.
@@ Appendix A Koopman operator and extended dynamic mode decomposition
${\bm{h}}_{t}=\Phi({\bm{h}}_{t})\,=\,(\underbrace{[\mathcal{K}\phi]\circ\cdots\circ[\mathcal{K}\phi]}_{t})({\bm{h}}_{0})=\sum_{k}\lambda^{t}_{k}\,\,\phi_{\lambda_{k}}({\bm{h}}_{0})\,\,{\bm{c}}^{\Phi}_{k}.$ (10)
@@ A.1 Extended dynamic mode decomposition
The core concept of EDMD involves approximating the operator’s action on $\mathcal{F}\,=\,L^{2}(\mathcal{H},\mu_{h})$ by selecting a finite dimensional subspace ~$\widetilde{\mathcal{F}}_{M}\subset\mathcal{F}$. To define this subspace, we start by choosing a dictionary $\mathcal{F}_{M}\,=\,\{\psi_{i}:\mathcal{H}\rightarrow\mathbb{R}\in\mathcal{H}\mid i=1,\cdots,M\}$. We have
@@
$\mathcal{J}\,=\,\frac{1}{2}\sum_{n=1}^{N}\left\lVert\rho({\bm{h}}_{n})\right\rVert^{2}=\frac{1}{2}\sum_{n=1}^{N}\left\lVert{\bm{a}}^{\mkern-1.5mu\mathsf{T}}\left(\mathcal{F}_{M}({\bm{h}}_{n}^{\prime})-K\mathcal{F}_{M}({\bm{h}}_{n})\right)\right\rVert^{2}.$ (11)
@@
where $\Lambda$ is a diagonal matrix with the eigenvalues. Due to numerical issues, one typically predicts one time step at the time, and projects down to state space each time.
In many applications, one is not necessarily interested in the spectral analysis, but only prediction. One then typically sets $\mathbb{F}=\mathbb{R}$, solves for $K$ in exact way, but solves for $C$ as
@@
For prediction one simply applies $K$ several times,
@@
In practice, one usually predicts step by step ${\bm{h}}_{t}$, that is, maps it down to the state space and maps back to the observable image space, before applying $K$ again. It is also worth noting that one may be more interested in mapping to an output of a function $f\in\mathcal{F}$ instead, and then one simply swaps $\Phi$ with $f$ when approximating $C$.
@@ A.2 Controlled dynamical systems and the Koopman operator
When approximating the Koopman operator for controlled systems with EDMD, the dictionary we choose needs alteration due to the domain ×$\mathcal{H}\times\ell(\mathcal{X})$ being infinite dimensional. Korda and Mezić [2018a] propose dictionaries that are both computable and enforce the linearity relationship assumed in Equation 13. The dictionaries to be considered are on the form
@@ Appendix B Theory
This restriction to the type of state and input spaces we consider is sufficient to construct all neural networks of interest by choosing pair of points from the space in question, and constructing the weight and bias as in Equation 4. It is also argued in Bolager et al. [2023] that most interesting real-world application will contain some noise and make sure that the state and input spaces is approximately on the form given in the definition.
@@ B.1 Uncontrolled systems
The following assumptions are made for the dictionary and the underlying system $F$:
  1. 1.
@@
then Item 3.1 holds. In particular, it holds when $\{\psi_{i}\}_{m=1}^{M}$ are independent tanh functions.
@@
to have measure zero, $\lambda(A)=0$, due to Proposition 1 by Mityagin [2020]. We have $\lambda(A\cap\mathcal{H})\leq\lambda(A)=0$. Finally due to absolutely continuous measure $\mu_{h}$, we have
@@
The following lemma makes sure we can circumvent assumptions made in Korda and Mezić [2018b], which requires on the dictionary in the EDMD algorithm to be an orthonormal basis (o.n.b.) of $\mathcal{F}$.
@@ B.2 Controlled systems
Proving convergence for controlled systems gives different challenges. The results above cannot easily be shown for controlled systems. The reason being that the dictionary space one uses is not a basis for the observables in the controlled setting, with the dynamical system extended by the left-shift operator, and the simplification made for EDMD in controlled systems. The results above may be extended, but the EDMD will not converge to the Koopman operator, but rather to $P^{\mu}_{\infty}\mathcal{K}_{\mathcal{F}_{\infty}}$, where $P^{\mu}_{\infty}$ is the $L^{2}(\mu)$ projection onto the closure of the dictionary space [Korda and Mezić, 2018a]. However, results exists for the continuous controlled systems, with certain convergences for the generator. This is sadly not as strong as above, but an interesting path to connect RNNs/NeuralODEs to such theory [Nüske et al., 2023].
@@ Appendix E Reservoir Models: Echo State Networks
An ESN consists of a reservoir and a readout. The reservoir contains neurons which are randomly connected to inputs and are these are not trained. For a certain discrete time $t\in\mathbb{Z}_{\geq 0}$ denoting the inputs as ${\bm{h}}_{t}\in\mathbb{R}^{N_{h}}$ and output as ${\bm{y}}_{t}\in\mathbb{R}^{N_{y}}$, and the internal reservoir states as ${\bm{k}}_{t}\in\mathbb{R}^{N_{k}}$, the reservoir provides an update rule for the internal units as
@@
where $f^{out}$ is the output activation, $W^{out}\in\mathbb{R}^{N_{y}\times N_{y}}$ are output weights and $({\bm{h}}_{t+1},{\bm{k}}_{t+1},{\bm{y}}_{t})$ denotes the concatenation of ${\bm{h}}_{t+1}$, ${\bm{k}}_{t+1}$ and ${\bm{y}}_{t}$. In the readout the model learns the connections from the reservoir to the readout, for example via (regularized) regression. A so-called feedback connection allows for the readout values to be fed back into the reservoir, as shown in Equation 21, establishing a recurrent relation.
@@ G.1 Geometrical measure
The Kullback–Leibler divergence of two probability densities $p({\bm{x}})$ and $q({\bm{x}})$ is defined as
@@ H.1 Sampled RNN
The implementation of our gradient-free training module for sampled RNNs was done in Python since the key tools for the algorithm already exist as Python libraries. The algorithm requires the ability to sample weights and biases, thus we used the Python library swimnetworks by Bolager et al. [2023]. Furthermore, we were able to alleviate the approximation of the Koopman operator, in the uncontrolled as well as controlled setting using some functionalities from the Python library datafold by Lehmberg et al. [2020].
Sampled RNNs have only a few hyperparameters: the number of nodes in the hidden layer, the activation function of the hidden layer, and a cutoff for small singular values in the least-squares solver. We often refer to the singular value cutoff hyperparameter as the regularization rate. In the case of a sampled RNN with time delay, additional hyperparameters are the number of time delays and the number of PCA components, if used.
Sampled RNNs do not only have a low fit time but also a short hyperparameter tuning since there are only a few degrees of freedom. Thus, our newly proposed method offers efficiency beyond the short training time.

### v2

@@ 4 Computational experiments
We now discuss a series of experiments designed to illustrate the benefits and challenges of our construction approach. We compare our method to the state-of-the-art iterative gradient-based method called shPLRNN, which we explain in Appendix D. For the real-world weather dataset, we also compare our approach with a long short-term memory (LSTM) model. Furthermore, due to the similarities our method bears with reservoir models, we also compare with an established reservoir model, namely an echo state network (ESN), further explained in Appendix E. Details on the datasets can be found in Appendix F, hyperparameters for all models are given in Appendix H, the evaluation metrics are explained in Appendix G and a further comparison discussion as well as the hardware details are provided in Appendix H. In addition to the experiments found in this section, Appendix I contains two more real-world data experiments and Appendix J contains a comparison study where we performed tests to observe the effects each of SWIM and the Koopman operator has on performance. The code to reproduce the experiments from the paper, as well as an up-to-date code base, can be found at
@@
https://gitlab.com/felix.dietrich/swimrnn.
In Table 1, we list the quantitative results of the experiments without control. Each entry stems from five different runs, where the random seed is changed in order to ensure a more robust result. We give the mean over these five runs, as well as the minimum and maximum among them. The results from the experiment with control are discussed in Section 4.4 and show in Figure 5.
@@ 4.1 Simple ODEs: Van der Pol Oscillator
We consider the Van der Pol oscillator system for a simple illustration of our method. A sampled RNN with a $tanh$ activation and a single hidden layer of width 80 is used, and the prediction method follows Algorithm 3.
The model is evaluated on test data; the averaged error and training time are reported in Table 1. One trajectory from the test set is visualized in Figure 2. It should be noted that predictions start with an initial condition from the test dataset which is used to make the first prediction, and afterwards continue using this prediction as an input to predict the next state, without information from the ground-truth dataset. In the results we observe a very stable trajectory over a long prediction horizon, and furthermore all eigenvalues of the Koopman operator are inside the unit disk (see Appendix C), thus we are certain that the model is stable. This experiment is also significant due to the periodic nature of the system, which is captured with our model, although neural network architectures in general struggle to capture periodicity (Ziyin et al., 2020). Compared to the gradient-descent trained shPLRNN our method is much faster and achieves higher forecasting accuracy, with lower MSE as prediction error. Compared with an ESN, our model has a shorter fit time and a smaller error. However, the hyperparameter search is simpler for our method since there are fewer hyperparameters to tune.
@@ 4.4 Example with control inputs: Forced Van der Pol Oscillator
We consider again the Van der Pol oscillator, where now the second coordinate, $h_{2}$, is controlled with an external input $x$, and using a sampled RNN model we perform model predictive control (MPC) as in Algorithm 4. We let $\mathcal{G}_{\hat{M}}$ be the identity function, and let $B$ map from $\mathcal{X}$ to $\mathbb{R}^{M}$ (adding nonlinearity for $x$ did not yield different results for this system, and is described in Section H.1.1). The sampled recurrent neural network is then the identity $\mathcal{G}_{\hat{M}}$ and $\mathcal{F}_{M}$ is a hidden layer of width 128 and $tanh$ activation. This network is then passed as a surrogate model to a linear-quadratic regulator (LQR). The network, in combination with the LQR, can successfully steer the state to the target state (see Figure 5). We consider five different runs, where only the random seed is varied, and obtain the mean controller cost to be 125.92 and the mean training time of 1.122 seconds. The norm of the state is also tracked over time, for five different runs we show the norms and the pointwise mean (over the runs) in Figure 5. This experiment highlights a key advantage of our model, which allows for modelling a nonlinear system such as the Van der Pol oscillator using a linear controller such as LQR. This implies that the well-established tools from linear control theory can be applied to non-linear systems using our method.
@@ 4.5 Example with real-world data
We apply our approach to the climate data presented in TensorFlow (2024). The dataset contains a time series of 14 weather parameters recorded in Jena (Germany) between January 1st, 2009, and December 31st, 2016. The data offers freedom in choosing the sizes of the time delay and prediction horizons. We decided to fix the time delay to one week and set two separate experiments with a prediction horizon of one day and one week. In the sampled RNN and LSTM experiments, we performed a grid search for each model with hyperparameters specified in Section H.4.1.
Table 1 shows the averaged training time and error metrics for the two selected horizons (day and week). We observe that the models perform similarly in the case of a shorter horizon, while sampling offers much faster training. When considering a one-week horizon, shPLRNN outperforms, while sampled RNN is still orders of magnitude faster. We also note the prediction horizon does not influence the training time of the sampled model that agrees with the Algorithm 2. When comparing predictions for the longer horizon in Figure 6, we notice that the LSTM struggles to predict the high-frequency fluctuations of the measurement, but the sampled RNN and shPLRNN successfully capture them. Figure 6 also highlights the deficiency of the MSE metric because a low mean error does not always correspond to accurate predictions, as illustrated by all three models. Overall, we conclude that sampled RNNs can successfully capture chaotic real-world dynamics and produce results comparable to the iterative models while offering a significant speed-up in training.
@@ 5 Conclusion
Benefits of the approach In many examples, we demonstrate that we can train accurate recurrent networks orders of magnitude faster compared to networks trained by iterative methods. We also observe that the training approach works with a very small amount of training data. The direct connection to Koopman operator theory allows us to draw on existing theoretical results for dynamical systems, which we use to prove convergence in the limit of infinite width.
@@
Remaining challenges and future work Recurrent networks are often used for tasks in computer vision and natural language processing. These tasks require network architectures beyond feed-forward networks, like convolutional neural networks or transformers, which means we currently struggle with high dimensional data. The sampling scheme we use to construct the hidden weights of the neurons is currently not useful in constructing parameters for such architectures, but it is an intriguing challenge to work towards sampling them. Remaining challenges in the theoretical work include extending the theory shown in this paper to controlled systems, as well as bridging the Koopman theory for continuous dynamical systems and NeuralODEs.
@@ Appendix A Koopman operator and extended dynamic mode decomposition
${\bm{h}}_{t}=\Phi({\bm{h}}_{t})\,=\,(\underbrace{[\mathcal{K}\circ\cdots\circ\mathcal{K}]}_{t}\phi)({\bm{h}}_{0})=\sum_{k}\lambda^{t}_{k}\,\,\phi_{\lambda_{k}}({\bm{h}}_{0})\,\,{\bm{c}}^{\Phi}_{k}.$ (10)
@@ A.1 Extended dynamic mode decomposition
The core concept of EDMD involves approximating the operator’s action on $\mathcal{F}\,=\,L^{2}(\mathcal{H},\mu_{h})$ by selecting a finite dimensional subspace ~$\widetilde{\mathcal{F}}_{M}\subset\mathcal{F}$. To define this subspace, we start by choosing a dictionary $\mathcal{F}_{M}\,=\,\{\psi_{i}:\mathcal{H}\rightarrow\mathbb{R}\in\mathcal{H}\mid i=1,\cdots,M\}$. We the have
@@
$\mathcal{J}\,=\,\frac{1}{2}\sum_{n=1}^{N}\left\lVert\rho({\bm{h}}_{n})\right\rVert^{2}=\frac{1}{2}\sum_{n=1}^{N}\left\lVert{\bm{a}}^{\mkern-1.5mu\mathsf{T}}\left(\mathcal{F}_{M}({\bm{h}}_{n}^{\prime})-K\mathcal{F}_{M}({\bm{h}}_{n})\right)\right\rVert^{2}$ (11)
@@
where $\Lambda$ is a diagonal matrix with the eigenvalues. Due to numerical issues, one typically predict one time step at the time, and project down to state space each time.
In many applications, one is not necessarily interested in the spectral analysis, but only prediction. One then typically set $\mathbb{F}=\mathbb{R}$, solve for $K$ in exact way, but solve for $C$ as
@@
For prediction one simply apply $K$ several times,
@@
In practice, one usually predict step by step ${\bm{h}}_{t}$, that is, maps it down to the state space and maps back to the observable image space, before applying $K$ again. It is also worth noting that one may be more interested in mapping to an output of a function $f\in\mathcal{F}$ instead, and then one simply swap $\Phi$ with $f$ when approximating $C$.
@@ A.2 Controlled dynamical systems and the Koopman operator
When approximating the Koopman operator for controlled systems with EDMD, the dictionary we choose needs alteration due to the domain ×$\mathcal{H}\times\ell(\mathcal{X})$ is infinite dimensional. Korda and Mezić [2018a] proposes dictionaries that are both computable and enforces the linearity relationship assumed in Equation 13. The dictionaries to be considered are on the form
@@ Appendix B Theory
This restriction to the type of state and input spaces we consider is sufficient to construct all neural networks of interest by choosing pair of points from the space in question, and construct the weight and bias as in Equation 4. It is also argued in Bolager et al. [2023] that most interesting real-world application will contain some noise and make sure that the state and input spaces is approximately on the form given in the definition.
@@ B.1 Uncontrolled systems
The following assumptions is made for the dictionary and the underlying system $F$:
  1. 3.1.
@@
then Item 3.1 is satisfied. In particular, it is satisfied when $\{\psi_{i}\}_{m=1}^{M}$ are independent tanh functions.
@@
to have measure zero, $\lambda(A)=0$, due to Proposition 1 in Mityagin [2020]. We have $\lambda(A\cap\mathcal{H})\leq\lambda(A)=0$. Finally due to absolutely continuous measure $\mu_{h}$, we have
@@
Following lemma makes sure we can circumvent assumptions made in Korda and Mezić [2018b], which requires on the dictionary in the EDMD algorithm to be an orthonormal basis (o.n.b.) of $\mathcal{F}$.
@@ B.2 Controlled systems
Proving convergence for controlled systems gives different challenges. The results above cannot easily be shown for controlled systems. The reason being that the dictionary space one use is not a basis for the observables in the controlled setting, with the dynamical system extended by the left-shift operator, and the simplification made for EDMD in controlled systems. The results above may be extended, but the EDMD will not converge to the Koopman operator, but rather to $P^{\mu}_{\infty}\mathcal{K}_{\mathcal{F}_{\infty}}$, where $P^{\mu}_{\infty}$ is the $L^{2}(\mu)$ projection onto the closure of the dictionary space [Korda and Mezić, 2018a]. However, results exists for the continuous controlled systems, with certain convergences for the generator. This is sadly not as strong as above, but an interesting path to connect RNNs/NeuralODEs to such theory [Nüske et al., 2023].
@@ Appendix E Reservoir Models: Echo State Networks
An ESN consists of a reservoir and a readout. The reservoir contains neurons which are randomly connected to inputs and are these are not trained. Denoting the inputs as ${\bm{h}}_{t}\in\mathbb{R}^{N_{h}}$ and output as ${\bm{y}}_{t}\in\mathbb{R}^{N_{y}}$, and the internal reservoir states as ${\bm{k}}_{t}\in\mathbb{R}^{N_{k}}$, the reservoir provides an update rule for the internal units as
@@
where $f^{out}$ is the output activation, $W^{out}\in\mathbb{R}^{N_{y}\times N_{y}}$ are output weights and $({\bm{h}}_{t+1},{\bm{k}}_{t+1},{\bm{y}}_{t})$ denotes the concatenation of ${\bm{h}}_{t+1}$, ${\bm{k}}_{t+1}$ and ${\bm{y}}_{t}$.
In the readout the model learns the connections from the reservoir to the readout, for example via (regularized) regression. A so-called feedback connection allows for the readout values to be fed back into the reservoir, as shown in Equation 21, establishing a recurrent relation.
@@ G.1 Geometrical measure
The Kullback–Leibler divergence of two probabiliti densities $p({\bm{x}})$ and $q({\bm{x}})$ is defined as
@@ H.1 Sampled RNN
The implementation of our gradient-free training module for sampled RNNs was done in Python since the key tools for the algorithm already exist as Python libraries.
The algorithm requires the ability to sample weights and biases, thus we used the Python library swimnetworks by Bolager et al. [2023]. Furthermore, we were able to alleviate the approximation of the Koopman operator, in the uncontrolled as well as controlled setting using some functionalities from the Python library datafold by Lehmberg et al. [2020]. Sampled RNNs have only a few hyperparameters: the number of nodes in the hidden layer, the activation function of the hidden layer, and a cutoff for small singular values in the least-squares solver. We often refer to the singular value cutoff hyperparameter as the regularization rate.
In the case of a sampled RNN with time delay, additional hyperparameters are the number of time delays and the number of PCA components, if used. Sampled RNNs do not only have a low fit time but also a short hyperparameter tuning since there are only a few degrees of freedom. Thus, our newly proposed method offers efficiency beyond the short training time.
@@
We also considered how the number of hidden units effects the model’s performance. For the number of nodes considered powers of two, specifically $2^{2}$ until $2^{9}$ for both chaotic models. Each run was repeated five times to provide errorbars from min to max in addition to the average value, show in Figure 11. We notice the typical trend of a decreasing error, with possible overfitting in the case of Rossler.
(a) Lorenz
(b) Rossler
Figure 11: Ablation of hidden units for chaotic systems.